OpenStudy (anonymous):

How would you solve this? (cos a- cosb)^2 +(sina-sinb)^2 = (cos(a-b)-1)^2+ (sin(a-b)-0)^2

OpenStudy (solomonzelman):

By expanding everything. You might need these identities. in(A+B)=sin A cos B + cos A sin B sin(A-B)=sin A cos B - cos A sin B cos(A+B)=cos A cos B - sin A sin B cos(A-B)=cos A cos B + sin A sin B

OpenStudy (solomonzelman):

\[Cos^2x+Sin^2x=1\]\[(a-b)^2=a^2-2ab+b^2\]

OpenStudy (solomonzelman):

Theese 6 is all what you need to know I think. You should be able to do this, (sorry)

OpenStudy (anonymous):

thankyou

OpenStudy (solomonzelman):

I didn't help you much, this is too much work...

OpenStudy (anonymous):

it's fine ty anyways

OpenStudy (solomonzelman):

Alright, then yw.

OpenStudy (solomonzelman):

ty for the medal... guess best would be trying to do it, I'll be guiding if you need more assistance.